OpenStudy (anonymous):

find the solution set by factoring and using zero product rule or by using the quadratic formula 2x^2+6x=2

OpenStudy (radar):

I would rearrange the equation as:\[2x ^{2}+6x-2=0\] then divide all terms by 2.\[x ^{2}+3x-1\] then use quadratic formula

OpenStudy (radar):

\[-3\pm \sqrt{9+4}\over2\]\[-3\pm \sqrt{13}\over2\]